If there are 2 red balls, 4 blue balls, 6 white balls and a certain number of black balls in a container and the odds of selecting a black ball are 1 out of 5, how many black balls are there?
Answer by Boreal(15235)

12+b=number of balls
get 1 choice
b/(b+12)=(1/5)
cross-multiply
5b=b+12
4b=12
b=3 black balls
3/15=1/5